# Machine learning in top quark physics at ATLAS and CMS

### v1

@@ Abstract
This note presents an overview of machine-learning-based techniques used in the study of the top quark. The research community has developed a diverse set of ideas and tools, including algorithms for the efficient reconstruction of recorded collision events and innovative methods for statistical inference. Recent applications by the ATLAS and CMS collaborations are also highlighted.
@@ 1 Introduction
This note reviews a collection of state-of-the-art ML algorithms addressing various aspects of top quark research, including top quark and event reconstruction, analysis strategies, and novel methods for statistical inference. An outlook on the impact of ML on future research at the upcoming HL-LHC is also provided.
@@ 2 Top quark reconstruction
The reconstruction of top quark events typically involves two stages. First, the inference of the direction(s) of undetected neutrino(s) from leptonically decaying top quarks, $\textrm{t}\to\textrm{b}\,\ell\nu$, if present. Second, the association of the remaining decay products (leptons and jets) to each top quark in the event. Various ML algorithms have been developed to address both tasks.
The $\nu$-flow approach [9, 10] utilizes a normalizing flow neural network (NN) conditioned on reconstructed event observables. This network maps the true neutrino direction vector, derived from the event generator truth, to a three-dimensional normal distribution. By sampling from this distribution, the likelihood of possible neutrino directions can be inferred from data, as illustrated in Fig. 1a for a representative event, demonstrating superior performance over regressing the values via a feed-forward NN or employing a W boson mass constraint.
@@
Figure 1: Inference of the pseudorapidity of the neutrino originating from a leptonically decaying top quark: (a) $\nu$-flow approach [9]; (b) spanet approach [11].
@@ 5 HL LHC outlook
The CMS collaboration has recently addressed the former challenge by introducing the dctr method [20, 21] for reweighting simulated samples. The study comprised the reweighting of samples to emulate parameter shifts used for estimating the impact by systematic uncertainties. An example variation involving the $h_{\textrm{damp}}$ parameter of the powheg event generator [22] is shown in Fig. 3a. The good agreement observed after reweighting demonstrates that this technique could replace the need for generating dedicated samples in the future.

### v2

@@ Abstract
This note presents an overview of current and potential future applications of machine-learning-based techniques in the study of the top quark. The research community has developed a diverse set of ideas and tools, including algorithms for the efficient reconstruction of recorded collision events and innovative methods for statistical inference. Recent applications of some techniques by the ATLAS and CMS collaborations are also highlighted.
@@ 1 Introduction
This note reviews a collection of state-of-the-art ML algorithms addressing various aspects of top quark research, including top quark and event reconstruction, analysis strategies, and novel methods for statistical inference, which could offer valuable contributions to the field in the future. An outlook on the impact of ML on future research at the upcoming HL-LHC is also provided.
@@ 2 Top quark reconstruction
[…]
The -flow approach [9, 10] utilizes a normalizing flow neural network (NN) conditioned on reconstructed event observables. This network maps the true neutrino direction vector, derived from the event generator truth, to a three-dimensional normal distribution. By sampling from this distribution, the likelihood of possible neutrino directions can be inferred from data, as illustrated in Fig. 1a for a representative event, illustrating superior performance over regressing the values via a feed-forward NN or employing a W boson mass constraint.
@@
Figure 1: Inference of the pseudorapidity of the neutrino originating from the leptonically decaying top quark in semileptonic $\mathrm{t}\bar{\mathrm{t}}$ events: (a) -flow approach [9]; (b) spanet approach [11].
@@ 5 HL LHC outlook
The CMS collaboration has recently addressed the former challenge by introducing the dctr method [20, 21] for reweighting simulated samples. The study comprised the reweighting of samples to emulate parameter shifts used for estimating the impact by systematic uncertainties. An example variation is shown in Fig. 3a involving the $h_{\textrm{damp}}$ parameter of the powheg event generator [22], which regulates the energy of additional radiations. Its effect on the sample cannot be described analytically but is approximated by the NN. The good agreement observed after reweighting demonstrates that this technique could replace the need for generating dedicated samples in the future, resulting in improved sustainability by skipping the computational needs of classical detector simulations.